OpenStudy (anonymous):

(cosΘ - cosΘ)^2 + (cos Θ + cos Θ)^2 1) sin^2Θ 2) 4 3) 8 4) cos^2Θ

zepdrix (zepdrix):

\[\Large\rm (\color{royalblue}{\cos \theta-\cos \theta})^2+(\cos \theta+\cos \theta)^2\]\[\Large\rm (\color{royalblue}{0})^2+(\cos \theta+\cos \theta)^2\]That... should simplify things down a bit, yah? 0_o

zepdrix (zepdrix):

With the other part as well,\[\Large\rm (\color{royalblue}{0})^2+(2\cos \theta)^2\]

zepdrix (zepdrix):

Which isn't going to line up with any of these options... hmmmm

random231 (random231):

what is the point of giving (cos(theta)-cos(theta))^2 when everyone knows its zero. Emmy will you please check the question again?

OpenStudy (anonymous):

I truly don't know i'm taking an exam and that's exactly what the question says

OpenStudy (freckles):

"(cosΘ - cosΘ)^2 + (cos Θ + cos Θ)^2"<---by you I wonder if you meant to write \[(\cos(\theta)-\sin(\theta))^2+(\cos(\theta)+\sin(\theta))^2\] but that would still be a choice given

OpenStudy (freckles):

not be a choice given*